OpenStudy (anonymous):

Recall that P is a projection matrix if and only if PT=P and P2=P. Also, recall that R is a reflection matrix if and only if RT=R and R2=I. Finally, recall that A is an orthogonal matrix if and only if AAT=I=ATA. Let R be a reflection matrix. Determine whether each of the following statements is true or false. 1. If R is invertible, then R=I. 2. I−R is a projection matrix. 3. R−I is a projection matrix. 4. det(R)=1 or det(R)=−1. 5. If S is also a projection matrix, then so is R+S. 6. 1 is an eigenvalue for R.

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

Can you clarify what P2=P means? Also, what R2=I means? By these, do you mean: \(P^2=P\) and \(R^2=I\)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Well, look at \(1\). Can you derive that from your given statements? If so, it's true. If not, it's false.

OpenStudy (anonymous):

it is false

OpenStudy (anonymous):

Correct. Do you know why?

OpenStudy (anonymous):

cus we can find many reflection matrices and they are not the identity

OpenStudy (anonymous):

That's a great way of looking at it. Mine was that if we have \(R^2=I\), then multiplying by the inverse (we can, since \(R\) is invertible), we have \(R=R^{-1}\). Question: Does \(P^{T}\) represent the transpose of \(P\)?

OpenStudy (anonymous):

Rephrase that: Are you talking about \(P^{T}\) when you say PT?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

second one is also false cus (R-I)^2 is not equal to (R-I)

OpenStudy (anonymous):

I'm not so certain. I'm still thinking on that one and the other two.

OpenStudy (anonymous):

Oops, the others, I should say. :p

OpenStudy (anonymous):

\[(I-R) \text{ is a projection matrix} \Rightarrow (I-R)^{T}=I-R\] Is the second part true? You can try to figure it out for yourself by working out general matrices.

OpenStudy (anonymous):

2 is false because the send condition of projection matrices says P^2=P

OpenStudy (anonymous):

\[ \begin{bmatrix} 1 & 0 & \dots & 0\\ 0 & 1 & \dots &0\\ \vdots & \vdots & \ddots &\vdots\\ 0 & 0 &\dots & 1 \end{bmatrix} - \begin{bmatrix} x_{11} & x_{21} & \dots & x_{n1}\\ x_{12} & x_{22} & \dots & x_{n2}\\ \vdots & \vdots & \ddots & \vdots\\ x_{1n} & x_{2n} & \dots & x_{nn} \end{bmatrix} = \begin{bmatrix} 1-x_{11} & x_{21} & \dots & x_{n1}\\ x_{12} & 1-x_{22} & \dots & x_{n2}\\ \vdots & \vdots & \ddots & \vdots\\ x_{1n} & x_{2n} & \dots & 1-x_{nn} \end{bmatrix} \] I'm not seeing how this could be equal to the transpose, so I see 2 as false, too.

OpenStudy (anonymous):

Hmm.. Wait...

OpenStudy (anonymous):

Okay. #2 is true. Do you know why?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Start working out what \(I-R\) is. Then create the transpose. Employing the properties of \(R\), you can thus show that \(I-R=(I-R)^{T}\). Give it a shot.

OpenStudy (anonymous):

however to complete that u should prove that (I-R)^2=(I-R)

OpenStudy (anonymous):

Hmm, good point.

OpenStudy (anonymous):

Yeah, I had forgot that condition. So I suppose it is false. I cannot see any way in which \((I-R)^2=(I-R)\).

OpenStudy (anonymous):

I think 3 is also false for nearly the same reasons. What do you think?

OpenStudy (anonymous):

yes it is false

OpenStudy (anonymous):

@eliassaab